Guaranteed investment certificate TAXATION

Greg invested $100,000 in a three-year guaranteed investment certificate (GIC) on September 1, Year 1. The GIC pays interest of 1% in Years 1 to 3. The interest compounds and is reinvested each year on the anniversary date of the GIC. The full amount —principal and all compounded interest —will be received upon maturity on August 31, Year 4.


Determine the minimum amount of income that Irene must report for Year 1 through Year 4 in respect of this investment.
Answer #1

The guaranteed investment contract is an obligation on the insurance company to pay a fixed rate of interest to the holder of this contract until maturity. Increase an interest rate will increase the burden on the company


Here Peter /Irene has invested $100000 in GIC on Sept 1 , Yr 1 @ 1% compounded annually ( when nothing is specifically given) . To find its matured amount on end of Yr 4 , we simply need Future Value formula,
FV = PV * ( 1+i)^n

PV = Present value = $100000

i = interest rate = 1%

no of yrs .(yr 1 sept to Yr 4 , Aug 31) =3 yrs
FV = $100000 * (1.01) ^3 = $100000 * 1.0303 = $103030
The minimum amount of income that Irene must report for Year 1 through Year 4 in respect of this investment
= 103030 - 100000 = 3030 ( Interest Income)


answered by: flower